What will be the value of molality for an aqueous solution of 10% w/w NaOH?

Grade:12th pass

1 Answers

Jayant Kishore
8 years ago
TECHNICAL ERROR , I’ve posted two times but it does’t appear
10 % w/w means 100g NaOH in 1000g solution 
Moles of solute = 100/40 = 2.5
Mass of solvent= 1000 – 100 = 900 g = 0.9 kg
molality = 2.5 / 0.9 = 2.77 m